# LiDARCrafter: Dynamic 4D World Modeling from LiDAR Sequences

### v1

@@ Abstract
[…]
Project Pagehttps://lidarcrafter.github.io
[…]
GitHubhttps://github.com/lidarcrafter/toolkit
@@ 1 Introduction
Generative world models are rapidly advancing the synthesis and interpretation of large-scale sensor data for autonomous driving (Hu et al. 2023; Mei et al. 2024). Recent work has concentrated on structured modalities such as video and occupancy grids, whose dense, regular structure fits image and voxel pipelines (Gao et al. 2023; Wang et al. 2024a). LiDAR, though indispensable for metric 3D geometry and all-weather perception, remains underexplored. The point clouds acquired by LiDAR sensors are often sparse, unordered, and irregular (Kong et al. 2023b; Liang et al. 2025; Liu et al. 2023; Xu et al. 2024). Therefore, techniques designed for images or grids do not transfer cleanly.
@@
Early efforts, such as LiDARGen, project full $360^{\circ}$ scans to range images and borrow pixel-based methods (Zyrianov, Zhu, and Wang 2022). Later approaches improve single-frame fidelity but stop short of dynamic sequences (Nakashima and Kurazume 2024; Ran, Guizilini, and Wang 2024). Multimodal systems like UniScene and GENESIS rely on occupancy or video as intermediaries, limiting LiDAR independence and increasing computation (Li et al. 2025; Guo et al. 2025). A dedicated 4D LiDAR world model is therefore still absent.
@@
Bridging this gap requires progress on three fronts. First, user controllability. Text prompts are the most accessible interface, yet they lack the spatial detail to specify LiDAR scenes, whereas precise inputs such as 3D boxes, HD maps, or trajectories demand costly annotation (Hu et al. 2023; Bian et al. 2025). Second, temporal consistency. A single frame cannot reveal occlusion patterns or object kinematics for downstream utilization, necessitating a full 4D generation model. Conventional cross-frame attention strategies neglect the geometric continuity inherent in point clouds. Third, standardized evaluation. Existing video world models benefit from established benchmarks that score visual quality, perception accuracy, and closed-loop applicability (Huang et al. 2024). LiDAR still lacks standard protocols for evaluating fidelity and consistency across perspectives.
@@
To close these gaps, we introduce LiDARCrafter, a unified framework for controllable 4D LiDAR sequence generation. The key is an explicit, object-centric 4D layout that captures geometry and motion while achieving accessible yet precise control. In Text2Layout, a large language model (LLM) lifts the user prompt to an ego-centric scene graph, and a tri-branch diffusion network predicts object boxes, trajectories, and shape priors, yielding the full 4D layout. Layout2Scene feeds the layout to a range-image diffusion network and produces a high-fidelity first scan. The explicit layout modeling also enables fine-grained control like insertion, deletion, and dragging. Scene2Seq then generates the remaining frames autoregressively, warping past points with motion priors to keep temporal coherence. To standardize 4D LiDAR world models assessment, we release an evaluation suite that scores scene, object, and sequence quality.
@@
Experiments on the nuScenes dataset (Caesar et al. 2020) demonstrate that our method achieves top performance in both single-frame fidelity and sequence coherence, while offering intuitive control and editing capabilities. This work establishes a new benchmark for LiDAR-based dynamic world modeling in autonomous driving.
@@
  • We present LiDARCrafter, the first 4D generative world model dedicated to LiDAR data, with superior controllability and spatiotemporal consistency.
  • We introduce a tri-branch 4D layout conditioned pipeline that turns language into an editable 4D layout and uses it to guide temporally stable LiDAR synthesis.
@@
Figure 2: Framework of LiDARCrafter. In the Text2Layout stage (cf. Sec. 3.1), the natural-language instruction is parsed into an ego-centric scene graph, and a tri-branch diffusion network generates 4D conditions for bounding boxes, future trajectories, and object point clouds. In the Layout2Scene stage (cf. Sec. 3.2), a range-image diffusion model uses these conditions to generate a static LiDAR frame. In the Scene2Seq stage (cf. Sec. 3.3), an autoregressive module warps historical points with ego and object motion priors to generate each subsequent frame, producing a temporally coherent LiDAR sequence.
@@ 2 Related Work
Driving Generative World Models. Generative world models aim to simulate environmental dynamics to aid autonomous driving decisions. Recent works primarily use video and occupancy representations. Video-based methods like GAIA-1 (Hu et al. 2023), Dreamforge (Mei et al. 2024), and Epona (Zhang et al. 2025) employ autoregressive modeling and richer conditioning for improved synthesis. MagicDrive (Gao et al. 2023) utilizes a bird’s-eye view (BEV) for temporal consistency (Wang et al. 2024b; Zhao et al. 2025; Ni et al. 2025a). Occupancy-based methods, such as OccWorld (Zheng et al. 2024a), GaussianWorld (Zuo et al. 2025b), QuadricFormer (Zuo et al. 2025a), OccSora (Wang et al. 2024a), and DynamicCity (Bian et al. 2025), provide structured spatial representations beneficial for downstream tasks. Multimodal models like UniScene (Li et al. 2025) and GENESIS (Guo et al. 2025) align cross-modal features for consistency. Despite its importance, LiDAR remains underexplored, mostly limited to forecasting-oriented methods (Liu, Zhao, and Rhinehart 2025; Zhang et al. 2023a).
LiDAR Point Cloud Generation. Early LiDAR generative methods, including LiDARGen (Zyrianov, Zhu, and Wang 2022), projected LiDAR point clouds into range images. Subsequent approaches like RangeLDM (Hu, Zhang, and Hu 2024), Text2LiDAR (Wu et al. 2024), and WeatherGen (Wu et al. 2025) introduced latent diffusion models (Rombach et al. 2022) or diverse conditioning signals. R2DM (Nakashima and Kurazume 2024) and R2Flow (Nakashima et al. 2024) adopted single-stage diffusion models (Ho, Jain, and Abbeel 2020) for geometric precision. UltraLiDAR (Xiong et al. 2023) utilized BEV-based VQ-VAE (Van Den Oord, Vinyals et al. 2017) for rich scene editing, further extended by OpenDWM (Ni et al. 2025b). X-Drive (Xie et al. 2024) and UniScene (Li et al. 2025) leveraged cross-modal consistency. Yet, dynamic sequence generation and fine-grained controllability remain unresolved.
Controllability in Scene Synthesis. Fine-grained control is crucial in generative modeling. Current methods depend on detailed inputs such as BEV semantic maps (Gao et al. 2023; Mei et al. 2024), HD maps (Swerdlow, Xu, and Zhou 2024), and 3D bounding boxes (Zhang et al. 2024; Yang et al. 2024), but preparing these inputs is labor-intensive. Text-based methods (Wu et al. 2024; Hu et al. 2023; Mei et al. 2024) offer simpler control but typically lack spatial precision. Two-stage indoor scene synthesis methods (Zhai et al. 2024; Yang et al. 2025b) use intermediate scene graphs to improve spatial control, recently extended to outdoor static occupancy grids (Liu et al. 2025; Yang et al. 2025a). Mature methods for controllable generation and editing of dynamic 4D LiDAR scenes, however, are not yet established.
@@ 3 LiDARCrafter: 4D LiDAR World Model
The cornerstone is an explicit 4D foreground layout that bridges the descriptive power of language and the geometric rigor required by LiDAR. As shown in Fig. 2, our framework adopts a three-stage process. In the Text2Layout stage (cf. Sec. 3.1), an LLM converts the instruction into an ego-centric scene graph, and a tri-branch diffusion sampler generates object boxes, trajectories, and shape priors, which serve as the conditioning layout signal. In the Layout2Scene stage (cf. Sec. 3.2), a range-image diffusion model turns the layout into a high-fidelity static scan. In the Scene2Seq stage (cf. Sec. 3.2), the static cloud is autoregressively warped and inpainted to yield drift-free frames. Finally, our EvalSuite (cf. Sec. 3.4) adds metrics for object semantics, layout soundness, and motion fidelity, giving the first comprehensive benchmark for 4D LiDAR generation.
@@ 3.1 Text2Layout: 4D Layout Generation
Natural-language prompts alone lack the spatial precision needed for complex world modeling. We therefore introduce a scene graph as an intermediate, explicit encoding of object geometry and relations. LLMs can parse text into such graphs, a strategy proven effective for scene synthesis (Feng et al. 2023; Yang et al. 2025b). LiDARCrafter extends this idea to dynamic outdoor settings. The LLM first builds a 4D scene graph from the prompt. A diffusion decoder then transforms this graph into a detailed layout of object boxes, trajectories, and shape priors that guides the subsequent LiDAR sequence generation process.
Language-Driven Graph Construction. Given a textual instruction, we prompt an LLM (Achiam et al. 2023) to build an ego-centric scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. A tailored query enumerates all foreground objects, producing the node set $\mathcal{V}=\{v_{0},\dots,v_{M}\}$, where $v_{0}$ denotes the ego vehicle and the remaining $M$ nodes represent dynamic objects. Each node $v_{i}$ is annotated with its semantic class $c_{i}$ and a motion state phrase $s_{i}$ (e.g., “go straight”). For every ordered pair $(i,j)$ with $i\neq j$, a directed edge $e_{i\to j}\in\mathcal{E}$ encodes their spatial relation (e.g., “in front of”, “lager than”, details in Fig. 2). Unlike prior work (Liu et al. 2025), including the ego node yields a structurally complete scene graph that fully conditions downstream layout generation.
@@
Graph-Fusion Encoder. To obtain context-aware priors for every tuple, following the method in the indoor area (Zhai et al. 2024), we process the scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with an $L$-layer TripletGCN (Johnson, Gupta, and Fei-Fei 2018). We first embed nodes and edges with a frozen CLIP text encoder (Radford et al. 2021) to bring richer semantics:
$\mathbf{h}_{v_{i}}^{(0)}=\texttt{concat}(\text{CLIP}(c_{i}),\,\text{CLIP}(s_{i}),\,\boldsymbol{\omega}_{i})$ (1)
@@
$({\tilde{\mathbf{h}}}_{v_{i}}^{(\ell)},\mathbf{h}_{e_{i\to j}}^{(\ell+1)},\tilde{\mathbf{h}}_{v_{j}}^{(\ell)})=\Phi_{\text{edge}}(\mathbf{h}_{v_{i}}^{(\ell)},\mathbf{h}_{e_{i\to j}}^{(\ell)},\mathbf{h}_{v_{j}}^{(\ell)})$ (2)
$\mathbf{h}_{v_{i}}^{(\ell+1)}=\tilde{\mathbf{h}}_{v_{i}}^{(\ell)}+\Phi_{\text{agg}}\Big{(}\texttt{avg}\big{(}\tilde{\mathbf{h}}_{{v}_{j}}^{(\ell)}~|~v_{j}\in N_{\mathcal{G}}(v_{i})\big{)}\Big{)}.$
@@
Layout Diffusion Decoder. The final node embeddings condition a tri-branch diffusion decoder (Rombach et al. 2022), one branch per element of $\mathcal{O}_{i}$. Let $\mathbf{d}^{o}_{\tau}$ be the noisy sample of modality $o\in\mathcal{O}_{i}$ at timestep $\tau$. Each branch minimizes:
@@
sharing a common noise schedule. Boxes and trajectories are denoised using a lightweight 1D U-Net (Ronneberger, Fischer, and Brox 2015), while object shapes are synthesised with a point-based U-Net (Zheng et al. 2024b). Unlike LOGen (Yan et al. 2025), we match only the LiDAR sampling distribution, not the exact foreground points, which eliminates the heavy DiT cost (Peebles and Xie 2023) yet still delivers plausible inputs for later refinement. More details about the denoiser are given in the Appendix.
Figure 3: Details of our range-image diffusion model.
@@ 3.2 Layout2Scene: Controlled LiDAR Generation
LiDARCrafter ensures generation fidelity by using a unified range-image diffusion backbone that generates LiDAR point clouds end to end. Given the scene graph $\mathcal{G}$ and the decoded layout $\mathcal{O}_{i}$, the network denoises Gaussian noise into the clean range frame $\mathbf{I}^{0}$, thereby bootstrapping the LiDAR sequence $\mathcal{P}=\{\mathbf{P}^{t}\}_{t=0}^{T}$ while following the pose, trajectory, and coarse shape of each object. The range-view representation preserves native LiDAR geometry while remaining convolution-friendly (Nakashima and Kurazume 2024; Kong et al. 2023a; Xu et al. 2025; Kong et al. 2025).
Sparse Object Conditioning. Directly projecting all foreground points into the range image, as in OLiDM (Yan et al. 2025), inadequately represents small or distant objects (e.g., a car at 1515m may occupy only a few dozen pixels). To address this, we condition the model on sparse object representations that encode semantics, pose, and coarse shape, thereby enabling the model to hallucinate fine structure, as shown in Fig. 3. For each node, we aggregate its features
@@
where $\pi(\mathbf{b}_{i})$ is the 3D box projected to image coordinates, $\Phi_{\text{pos}}$ is a positional embedder, and $\Phi_{\text{cls}},\Phi_{\text{box}}$ are learned MLPs. A lightweight self-attention layer diffuses contextual cues across tokens (Vaswani et al. 2017), producing the refined vector ${\mathbf{h}}_{v_{i}}$. The ego token is further compressed by an MLP to form a scene-level vector $\mathbf{h}_{\text{ego}}$.
@@
Layout-Driven Scene Editing. As each frame is anchored by an explicit layout, we can edit objects without disturbing the static background, which is crucial for testing planners. After the original scene $\mathbf{d}_{0}^{\text{orig}}$ is synthesized, a user may alter the layout tuple. We then rerun the reverse diffusion, preserving pixels whose 2D projections remain unchanged following (Lugmayr et al. 2022). At each denoising step:

### v3

# ] [Uncaptioned image] WorldBench Team

@@ Abstract
[…]
[ [Uncaptioned image] Project Page]https://lidarcrafter.github.io
[…]
@@ 1 Introduction
Generative world models are rapidly advancing the synthesis and interpretation of large-scale sensor data for autonomous driving [18, 10, 28, 53]. Recent work has concentrated on structured modalities such as video and occupancy grids, whose dense, regular structure fits image and voxel pipelines [7, 44]. LiDAR, though indispensable for metric 3D geometry and all-weather perception, remains underexplored. The point clouds acquired by LiDAR sensors are often sparse, unordered, and irregular [16, 21, 23, 51, 48, 24, 20, 66]. Therefore, techniques designed for images or grids do not transfer cleanly.
@@
Early efforts, such as LiDARGen, project full $360^{\circ}$ scans to range images and borrow pixel-based methods [69]. Later approaches improve single-frame fidelity but stop short of dynamic sequences [30, 37]. Multimodal systems like UniScene and GENESIS rely on occupancy or video as intermediaries, limiting LiDAR independence and increasing computation [19, 8]. A dedicated 4D LiDAR world model is therefore still absent.
@@
Bridging this gap requires progress on three fronts. First, user controllability. Text prompts are the most accessible interface, yet they lack the spatial detail to specify LiDAR scenes, whereas precise inputs such as 3D boxes, HD maps, or trajectories demand costly annotation [10, 2]. Second, temporal consistency. A single frame cannot reveal occlusion patterns or object kinematics for downstream utilization, necessitating a full 4D generation model. Conventional cross-frame attention strategies neglect the geometric continuity inherent in point clouds. Third, standardized evaluation. Existing video world models benefit from established benchmarks that score visual quality, perception accuracy, and closed-loop applicability [12]. LiDAR still lacks standard protocols for evaluating fidelity and consistency across perspectives.
@@
To close these gaps, we introduce LiDARCrafter, a unified framework for controllable 4D LiDAR sequence generation. The key is an explicit, object-centric 4D layout that captures geometry and motion while achieving accessible yet precise control (see Figure 1). In Text2Layout, a large language model (LLM) lifts the user prompt to an ego-centric scene graph, and a tri-branch diffusion network predicts object boxes, trajectories, and shape priors, yielding the full 4D layout. Layout2Scene feeds the layout to a range-image diffusion network and produces a high-fidelity first scan. The explicit layout modeling also enables fine-grained control like insertion, deletion, and dragging. Scene2Seq then generates the remaining frames autoregressively, warping past points with motion priors to keep temporal coherence. To standardize 4D LiDAR world models assessment, we release an evaluation suite that scores scene, object, and sequence quality.
@@
Experiments on the nuScenes dataset [3] demonstrate that our method achieves top performance in both single-frame fidelity and sequence coherence, while offering intuitive control and editing capabilities. This work establishes a new benchmark for LiDAR-based dynamic world modeling in autonomous driving.
@@
  • We present LiDARCrafter, the first 4D generative world model dedicated to LiDAR data, with superior controllability, spatiotemporal consistency, and autoregressive generation capability.
  • We introduce a tri-branch 4D layout conditioned pipeline that turns language into an editable 4D layout and uses it to guide temporally stable LiDAR sequence synthesis.
@@
Figure 2: Framework of LiDARCrafter. In the Text2Layout stage (cf. Section 3.1), the natural-language instruction is parsed into an ego-centric scene graph, and a tri-branch diffusion network generates 4D conditions for bounding boxes, future trajectories, and object point clouds. In the Layout2Scene stage (cf. Section 3.2), a range-image diffusion model uses these conditions to generate a static LiDAR frame. In the Scene2Seq stage (cf. Section 3.3), an autoregressive module warps historical points with ego and object motion priors to generate each subsequent frame, producing a temporally coherent LiDAR sequence.
@@ 2 Related Work
Driving Generative World Models. Generative world models aim to simulate environmental dynamics to aid autonomous driving decisions. Recent works primarily use video and occupancy representations. Video-based methods like GAIA-1 [10], DreamForge [28], and Epona [60] employ autoregressive modeling and richer conditioning for improved synthesis. MagicDrive [7] utilizes a bird’s-eye view (BEV) for temporal consistency [45, 63, 32]. Occupancy-based methods, such as OccWorld [64], GaussianWorld [68], QuadricFormer [67], OccSora [44], and DynamicCity [2], provide structured spatial representations beneficial for downstream tasks. Multimodal models like UniScene [19] and GENESIS [8] align cross-modal features for consistency. Despite its importance, LiDAR remains underexplored, mostly limited to forecasting-oriented methods [22, 61].
LiDAR Point Cloud Generation. Early LiDAR generative methods, including LiDARGen [69], projected LiDAR point clouds into range images. Subsequent approaches like RangeLDM [11], Text2LiDAR [46], and WeatherGen [47] introduced latent diffusion models [38] or diverse conditioning signals. R2DM [30] and R2Flow [31] adopted single-stage diffusion models [9] for geometric precision. UltraLiDAR [50] utilized BEV-based VQ-VAE [42] for rich scene editing, further extended by OpenDWM [33]. X-Drive [49] and UniScene [19] leveraged cross-modal consistency. Yet, dynamic 4D sequence generation and fine-grained controllability remain an unresolved problem.
Controllability in Scene Synthesis. Fine-grained control is crucial in generative modeling. Current methods depend on detailed inputs such as BEV semantic maps [7, 28], HD maps [41], and 3D bounding boxes [59, 55], but preparing these inputs is labor-intensive. Text-based methods [46, 10, 28] offer simpler control but typically lack spatial precision. Two-stage indoor scene synthesis methods [58, 57] use intermediate scene graphs to improve spatial control, recently extended to outdoor static occupancy grids [25, 56]. Mature methods for controllable generation and editing of dynamic 4D LiDAR scenes, however, are not yet established.
@@ 3 LiDARCrafter: 4D LiDAR World Model
The cornerstone is an explicit 4D foreground layout that bridges the descriptive power of language and the geometric rigor required by LiDAR. As shown in Figure 2, our framework adopts a three-stage process. In the Text2Layout stage (cf. Section 3.1), an LLM converts the instruction into an ego-centric scene graph, and a tri-branch diffusion sampler generates object boxes, trajectories, and shape priors, which serve as the conditioning layout signal. In the Layout2Scene stage (cf. Section 3.2), a range-image diffusion model turns the layout into a high-fidelity static scan. In the Scene2Seq stage (cf. Section 3.2), the static cloud is autoregressively warped and inpainted to yield drift-free frames. Finally, our EvalSuite (cf. Section 3.4) adds metrics for object semantics, layout soundness, and motion fidelity, giving the first comprehensive benchmark for 4D LiDAR generation.
@@ 3.1 Text2Layout: 4D Layout Generation
Natural-language prompts alone lack the spatial precision needed for complex world modeling. We therefore introduce a scene graph as an intermediate, explicit encoding of object geometry and relations. LLMs can parse text into such graphs, a strategy proven effective for scene synthesis [6, 57]. LiDARCrafter extends this idea to dynamic outdoor settings. The LLM first builds a 4D scene graph from the prompt. A diffusion decoder then transforms this graph into a detailed layout of object boxes, trajectories, and shape priors that guides the subsequent LiDAR sequence generation process.
Language-Driven Graph Construction. Given a textual instruction, we prompt an LLM [1] to build an ego-centric scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. A tailored query enumerates all foreground objects, producing the node set $\mathcal{V}=\{v_{0},\dots,v_{M}\}$, where $v_{0}$ denotes the ego vehicle and the remaining $M$ nodes represent dynamic objects. Each node $v_{i}$ is annotated with its semantic class $c_{i}$ and a motion state phrase $s_{i}$ (e.g., “go straight”). For every ordered pair $(i,j)$ with $i\neq j$, a directed edge $e_{i\to j}\in\mathcal{E}$ encodes their spatial relation (e.g., “in front of”, “lager than”, details in Figure 2). Unlike prior work [25], including the ego node yields a structurally complete scene graph that fully conditions downstream layout generation.
@@
Graph-Fusion Encoder. To obtain context-aware priors for every tuple, following the method in the indoor area [58], we process the scene graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with an $L$-layer TripletGCN [13]. We first embed nodes and edges with a frozen CLIP text encoder [36] to bring richer semantics:
$\mathbf{h}_{v_{i}}^{(0)}=\texttt{concat}(\text{CLIP}(c_{i}),\,\text{CLIP}(s_{i}),\,\boldsymbol{\omega}_{i}),$ (1)
@@
$({\tilde{\mathbf{h}}}_{v_{i}}^{(\ell)},\mathbf{h}_{e_{i\to j}}^{(\ell+1)},\tilde{\mathbf{h}}_{v_{j}}^{(\ell)})=\Phi_{\text{edge}}(\mathbf{h}_{v_{i}}^{(\ell)},\mathbf{h}_{e_{i\to j}}^{(\ell)},\mathbf{h}_{v_{j}}^{(\ell)}),$ (2)
$\mathbf{h}_{v_{i}}^{(\ell+1)}=\tilde{\mathbf{h}}_{v_{i}}^{(\ell)}+\Phi_{\text{agg}}\Big(\texttt{avg}\big(\tilde{\mathbf{h}}_{{v}_{j}}^{(\ell)}\penalty 10000\ |\penalty 10000\ v_{j}\in N_{\mathcal{G}}(v_{i})\big)\Big).$
@@
Layout Diffusion Decoder. The final node embeddings condition a tri-branch diffusion decoder [38], one branch per element of $\mathcal{O}_{i}$. Let $\mathbf{d}^{o}_{\tau}$ be the noisy sample of modality $o\in\mathcal{O}_{i}$ at timestep $\tau$. Each branch minimizes:
@@
sharing a common noise schedule. Boxes and trajectories are denoised using a lightweight 1D U-Net [39], while object shapes are synthesised with a point-based U-Net [65]. Different from LOGen [14], we match only the LiDAR sampling distribution, not the exact foreground points, which eliminates the heavy DiT cost [34] yet still delivers plausible inputs for later refinement. More details about the denoiser are given in the Appendix.
@@ 3.2 Layout2Scene: Controlled LiDAR Generation
Figure 3: Structures of our range-image LiDAR diffusion model.
LiDARCrafter ensures generation fidelity by using a unified range-image diffusion backbone that generates LiDAR point clouds end to end. Given the scene graph $\mathcal{G}$ and the decoded layout $\mathcal{O}_{i}$, the network denoises Gaussian noise into the clean range frame $\mathbf{I}^{0}$, thereby bootstrapping the LiDAR sequence $\mathcal{P}=\{\mathbf{P}^{t}\}_{t=0}^{T}$ while following the pose, trajectory, and coarse shape of each object. The range-view representation preserves native LiDAR geometry while remaining convolution-friendly [30, 15, 52, 17].
Sparse Object Conditioning. Directly projecting all foreground points into the range image, as in OLiDM [54], inadequately represents small or distant objects (e.g., a car at 1515m may occupy only a few dozen pixels). To address this, we condition the model on sparse object representations that encode semantics, pose, and coarse shape, thereby enabling the model to hallucinate fine structure, as shown in Figure 3. For each node, we aggregate its features
@@
where $\pi(\mathbf{b}_{i})$ is the 3D box projected to image coordinates, $\Phi_{\text{pos}}$ is a positional embedder, and $\Phi_{\text{cls}},\Phi_{\text{box}}$ are learned MLPs. A lightweight self-attention layer diffuses contextual cues across tokens [43], producing the refined vector ${\mathbf{h}}_{v_{i}}$. The ego token is further compressed by an MLP to form a scene-level vector $\mathbf{h}_{\text{ego}}$.
@@
Layout-Driven Scene Editing. As each frame is anchored by an explicit layout, we can edit objects without disturbing the static background, which is crucial for testing planners. After the original scene $\mathbf{d}_{0}^{\text{orig}}$ is synthesized, a user may alter the layout tuple. We then rerun the reverse diffusion, preserving pixels whose 2D projections remain unchanged following [26]. At each denoising step: